The half-life of carbon-14 is 5700 years. Find the age of a sample at which 40% of the radioactive nuclei originally present have decayed
Answer by lwsshak3(11628)

The half-life of carbon-14 is 5700 years. Find the age of a sample at which 40% of the radioactive nuclei originally present have decayed
**
Carbon 14 Dating Formula:A=Ao2^-t/h, A=amt of radioactive material remaining after time t, Ao=amt initially present, h=half life of the material.
..
Amt of material present=60%
60%Ao=Ao*2^-t/5700
0.6=2^-t/5700
take log of both sides
log(.6)=-t/5700 log(2)
-t/5700=log(.6)/log(2)
-t=[log(.6)/log(2)]*5700≈-4200
sample is approximately 4200 years old